the length of a rectangle is 3 meters longer than the width. If the diagonal of the rectangle is 15 meters. Find the length and width
Answer by ReadingBoosters(3246)

l = 3+w
diagonal = 15m
Using pythagorean theorem, a^2+b^2=c^2
(3+w)^2 + w^2 = 15^2
w^2+6w+9+w^2=225
2w^2 + 6w - 216 = 0
2(w^2 + 3w - 108)=0
2(w-9)(w+12)=0
w-9=0 so, w=9
w+12=0, so w=-12 NO
using the width of 9m
length = 3+9 = 12m
...............................
Delighted to help
-Reading Boosters
We want for others what we want for ourselves.